A current through galvanometer #Ig# is 2mA give full scale deflection with #Rg# 15ohms the amount of shunt connected to it to convert it into 10A ammeter is a) 0.030 b) 3 c) 5?

Grade:12th pass

1 Answers

ashish
6 years ago
dear rabik  
as  Ig *Rg =(I-Ig)S  (S = SHUNT)
2*10-3 *15= (10-2*10-3) S
30*10-3 =10S (10-2*10-3 =10 approx)
3*10-3ohm =S  Ans